# Not All Data are Good Labels: On the Self-supervised Labeling for Time Series Forecasting

### v3

@@ 1 Introduction
Time Series Forecasting (TSF) is a crucial task with extensive applications in energy, finance, engineering, and many other domains. Recent advances in deep learning have resulted in TSF methods that outperform traditional methods in precision, robustness, and scalability [53, 31, 36].
Nevertheless, deep learning-based TSF methods still face significant challenges such as overfitting, dependence on high-quality datasets, and inconsistent performance across datasets — issues exacerbated by flawed evaluation practices [34, 30]. Central to these challenges is the problem of low-quality labels [^1] In TSF, given two adjacent windows in a time series, $x$ (input window) and $y$ (output window), “labels” refer to $y$ fitted by predictions $\hat{y}=f(x;\theta)$ in a supervised learning setting., associated with inherent noise and anomalies in raw data. Existing strategies, such as multimodal data integration [43, 4] and dataset scaling [35], focus on augmenting or refining datasets but fail to address the core limitation: the reliance on raw labels as ground truth. To address this, we pose two critical questions:
@@
To better analyze overfitting, we derive a mask form of the co-objective loss, which partitions original time series into distinct components (see Figure 1). We employ a criterion, namely the sharpness metric $\lambda_{max}$ [10], which detects overfitting by evaluating the sharpness of loss landscape. Using $\lambda_{max}$ and practical evaluations of the decomposed loss, we identify the overfitted components. This process, termed Self-Correction with Adaptive Mask (Scam), discards raw data labels ($y$) based on reconstruction results ($\tilde{y}$) and current predictions ($\hat{y}$), smoothing the loss landscape and enhancing the generalization of the predictor model $f(\cdot;\theta)$.
We apply masked updates on a reconstruction network $g(\cdot;\phi)$ and a predictor model $f(\cdot;\theta)$ during training. During inference, the prediction will be generated directly by the updated $f(\cdot;\theta)$ with no additional cost. To further generalize across models of varying complexities, we also propose to apply Spectral Norm Regularization [46, 26] to parameters of the first or last linear layers without altering optimizers like previous works [10].
@@ 2.2 Proposed $g(\cdot;\phi)$ for Reconstruction
As depicted in Figure 2, the proposed $g(\cdot;\phi)$ comprises four convolutional layers, a cross-layer concatenation for multi-resolution integration, and a lightweight feedforward network (FFN) to decode the reconstructed results. The convolutional layers primarily act as a parameterized smoothing mechanism, similar to techniques for seasonal-trend decomposition [48, 16, 17]. The FFN then mixes information from different positions in a time series to reconstruct each data point using features extracted by convolutional layers (further details in Appendix D). Without introducing extra noise or patch-wise/point-wise masks [27], we directly learn an identity mapping.
In essence, $g(\cdot;\phi)$ is designed to traverse diverse alternatives with varying levels of fidelity, which differs in purpose from the usual reconstruction task [22, 24, 18]. Due to this difference, the architecture of $g(\cdot;\phi)$ might benefit from novel designs; however, this is not examined in this study. The proposed $g(\cdot;\phi)$ is merely a simple prototype to validate our ideas and is not claimed superior to other unexplored options for this novel task.
@@ 3.3 Self-Correction with Adaptive Mask (Scam)
Further evidence is provided by analyzing the loss landscape. A sharpness metric for optimization, proposed by [10], measures the generalization capability of a model. Specifically, the sharpness metric is defined as $\lambda_{max}=\|\nabla^{2}_{\theta}\mathcal{L}\|_{2}^{2}$, the largest eigenvalue of the Hessian matrix. A higher $\lambda_{max}$ indicates a sharper loss landscape, which correlates with a worse generalization or more severe overfitting. By computing this metric on the converged parameters, we observe that $\ell_{target}\odot M$ does exhibit a sharper landscape compared to $\ell_{target}\odot\overline{M}$. To address this, we keep the less sharp part of the $\mathcal{L}_{sup}$ term, i.e., $\mathcal{L}_{sup}\odot\overline{M}$, resulting in the revised loss:
@@ 3.4 Spectral Norm Regularization (snr)
Note that we have omitted the gradient constraints in Eq. 3. This term is, in fact, positively correlated with $\nabla_{\theta}f(x;\theta)$ (discussed in Appendix E). This further supports our previous analysis, as we have only tested MLP models, which converge easily with lower $\nabla_{\theta}f$. However, when the predictor is replaced by a Transformer-based model, the dominant source of overfitting shifts from $\ell_{target}$ to $\ell_{pred}$. This phenomenon is not unique to our self-supervised learning paradigm. For example, [10] proposed sharpness-aware optimization to address overfitting in supervised settings. While gradient penalties are theoretically effective, they may not be practical for complex models like Transformers, as computing second-order derivatives can significantly hinder optimization.
@@
A more direct approach is to regularize parameters using the sharpness metric, known as Spectral Norm Regularization [26, 46]:
@@
where $\|\cdot\|_{2}$ is the spectral norm (the largest eigenvalue of parameter matrix) and $\gamma$ is a learnable scale factor. When applying to self-attention (SA) in Transformer-based architecture, snr significantly undermines the expressiveness of attention score matrices (entropy collapse [49]). Hence, Ilbert et al. [10] conclude that snr is inapplicable to SA parameters. However, we observe that linear layers — typically the embedding layer before SA and the projection layer after SA — also contribute to the overall sharpness of the loss landscape. Consequently, we propose applying snr selectively to the pre- and post-SA linear layers. In this way, snr can work with MLP-based models as they are also composed of multiple linear layers. Further empirical studies on snr are discussed in Section E.3.
@@ 4 Experiment and Analysis
Our main evaluation involves seven datasets: Electricity, Weather, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), all of which are well-established TSF benchmarks and publicly available [40]. We also test the proposals using four PeMS datasets of a larger scale, as reported in Appendix 7. The predictor (i.e., the backbone model integrated with Scam) covers representative TSF models, including both MLP-based and Transformer-based architectures. Mlp [15] is a vanilla 2-layer baseline equipped with RevIN [12] while cycleNet [17] is a SOTA MLP-based model explicitly capturing cyclic trend in time series. PatchTST [27] and iTransformer [21] are Transformer-based models, representing channel-independent and channel-dependent methods, respectively. Following previous settings [53, 40, 41] for direct and fair comparison, we set prediction length $H\in\{96,192,336,720\}$ and look-back length to 96 for all datasets. We provide dataset descriptions, implementation details, and reproduction instructions in Appendix G.
@@ 4.1 Main Experiment (Q1)
Regarding Q1, our method demonstrates general effectiveness across various backbones and datasets. A well-known discrepancy between MLP-based and Transformer-based models is their dataset preferences: Transformer-based methods excel on large, regular datasets, while MLP-based methods perform better on noisy datasets [30, 34]. Scam helps bridge this gap by enabling Transformer-based models to perform competitively on traditionally challenging datasets and enhancing the robustness of MLP-based models.
@@ 4.2 Ablation Study (Q2)
snr, a practical alternative to the gradient penalty in Eq. 8, consistently enhances performance across backbones. However, iTrans, being more prone to overfitting on small datasets (discussed in [10]), benefits less from snr compared to cycleNet, likely due to insufficient data for generalization in higher-complexity architectures. With Scam, both models exhibit significant improvements, leveraging the expanded effective training data.
@@ 4.3 Scam: A Multiple Instance Learning View (Q3)
Multiple Instance Learning (MIL [25, 3]) is a classical weakly-supervised binary classification problem where typically a bag of instances is labeled positive if at least one instance is positive. [8, 5] extend MIL to Time Series Classification by treating an input window as a bag of instances (time points), enabling a model to predict based on instance-level classifications that are more interpretable.
@@ 5 Related Work
Meta-Learning for Time Series. Meta-Learning, by definition, seeks to perform in a learn-to-learn paradigm. Generally speaking, meta-learning for TSF includes optimization-based and metric-based methods. Optimization-based methods target optimal initial parameters [9], often involving a two-loop optimization process [28, 38]. Metric-based methods [7, 38] learn a metric function that provides more expressive measurements of distances between data samples, commonly comparing training and validation samples.
Our method Scam aligns with the broader scope of meta-learning. Specifically, the grid search (Section 3.1, Appendix B) follows the two-loop structure similar to optimization-based methods. However, it diverges by focusing on dataset space rather than parameter space. Additionally, the final mask form of Scam in essence provides a more accurate metric tailored for a supervised-learning setting. Previous works [38, 7] learn metrics on the sample level (a window of time series), whereas ours focuses on the instance level (individual data points of time series).
Self-supervised Learning for Time Series. Self-supervised learning trains models without relying on manually labeled data by using auxiliary tasks like generation, contrast, and reconstruction to learn expressive representation or create pseudo labels. In the realm of time series, this approach is discussed more in Time Series Classification (TSC) [11, 42, 22]. Recent works [8, 5] present a novel perspective that instances in time series/segments can have multiple labels. They propose corresponding weakly-supervised-learning methods that significantly improve both performance and interpretability.
@@ Appendix A Limitations
Our work constrains the discussion for TSF models in MLP- and Transformer-based ones, excluding other potentially feasible solutions [23, 1]. We omit these newly explored architectures since they have not outperformed the baseline methods we have adopted under the same budget. Also, we have not verified our method on LLM-based models (or models of an LLM-level scale) [39, 2] due to the excessive costs required.
@@ Appendix C Further Discussion on Gradient Constraint
For readers who are familiar with Reinforcement Learning, this derivation resembles the transfer from a constrained optimization to a penalized one (e.g., from TRPO [32] to PPO [33]). In brief, while the penalized form is theoretically equivalent to the constrained form, it is challenging to choose a fixed $\beta$ that works universally across all datasets or even within a single dataset (because intrinsic characteristics can vary over time). Thus, a more general form of constraint is required to better serve the penalty, similar to the concept of gradient clipping in PPO.
@@
However, in practice, snr have limitations when applied to the parameter matrix in self-attention mechanisms. This phenomenon is termed entropy catastrophe as discussed by [10]. In this paper, by analyzing the sharpness of different components in the predictor model, we propose to use pre-snr and post-snr combined, which specifically normalizes the first and last linear layer in the TSF models (see Section 3.4).
@@ D.1 Conv-concat Layer
Effective Receptive Field. As what was proposed by [23], the Effective Receptive Field (ERF) is a reasonable consideration for designing convolution-based architectures. To evaluate the ERF of our conv-concat layer, we input an impulse function and visualize the resulting ERF, as shown in Figure 10. The visualization demonstrates that, without requiring an extra-large convolution kernel, our proposed method achieves a near-global ERF. This is made possible by combining the outputs from different layers, each capturing distinct frequency patterns.
@@ D.2 Point-wise FFN
  • Patch-dependent Design: PatchTST [27] adopts a patch-dependent linear projector (shown in Figure 11(a)) that first flattens all features within patches and utilizes an extra-large weight matrix of shape $Ld\times d$, where $L$ is the sequence length and $d$ is the dimension of the latent embedding.
  • Patch-independent Design: [14] proposes that the necessity for patch-dependent designs depends on the specific task. For instance, tasks like forecasting may require patch-dependent projectors, while tasks like contrastive learning might favor a patch-independent design (shown in Figure 11(b)).
@@ E.2 Demystifying the Self-supervision in Scam
Channel-wise Distribution Alignment. Distribution shift, a core challenge in long-term Time Series Forecasting (TSF), is decisive for a TSF model to generalize on future data after training. The most widely adopted approach to address this issue is Reversible Instance Normalization (RevIn) [12, 20], which aligns the distributions of historical and future data.
@@
For models designed with channel-independence (CI), such as PatchTST [27] and cycleNet [17], aligning distributions across channels is of critical importance, especially when trained on datasets with a large number of variates. The channel-wise alignment introduced by our self-supervised reconstruction task provides an effective solution to distribution shifts, thereby enhancing the performance of CI predictor models.
@@ E.3 In-depth Examination on snr
We further verify the effectiveness of snr by analyzing the loss landscape. Transformer-based models, as discussed, are more prone to overfitting, particularly on noisy datasets. This weakens our adaptive mask’s ability to discard overfitted components, as $\ell_{pred}$ — the predictor’s training loss — also contributes to overfitting. Gradient penalty offers a solution to this issue. For TSF models specifically, where both inputs and outputs are time series with abrupt distribution shifts, parameter robustness is critical for mitigating overfitting. Previous works [46, 26] have shown that regulating parameters via their spectral norm enhances stability against input perturbations, which is particularly beneficial in TSF.

### v4

@@ 1 Introduction
Time Series Forecasting (TSF) is a crucial task with extensive applications in energy, finance, engineering, and many other domains. Recent advances in deep learning have resulted in TSF methods that outperform traditional methods in precision, robustness, and scalability [informer, deepar, prophet].
Nevertheless, deep learning-based TSF methods still face significant challenges such as overfitting, dependence on high-quality datasets, and inconsistent performance across datasets — issues exacerbated by flawed evaluation practices [benchmark_tkde, tfb]. Central to these challenges is the problem of low-quality labels [^1] In TSF, given two adjacent windows in a time series, $x$ (input window) and $y$ (output window), “labels” refer to $y$ fitted by predictions $\hat{y}=f(x;\theta)$ in a supervised learning setting., associated with inherent noise and anomalies in raw data. Existing strategies, such as multimodal data integration [textcues, visionts] and dataset scaling [tscaling], focus on augmenting or refining datasets but fail to address the core limitation: the reliance on raw labels as ground truth. To address this, we pose two critical questions:
@@
To better analyze overfitting, we derive a mask form of the co-objective loss, which partitions original time series into distinct components (see Figure 1). We employ a criterion, namely the sharpness metric $\lambda_{max}$ [samformer], which detects overfitting by evaluating the sharpness of loss landscape. Using $\lambda_{max}$ and practical evaluations of the decomposed loss, we identify the overfitted components. This process, termed Self-Correction with Adaptive Mask (Scam), discards raw data labels ($y$) based on reconstruction results ($\tilde{y}$) and current predictions ($\hat{y}$), smoothing the loss landscape and enhancing the generalization of the predictor model $f(\cdot;\theta)$.
We apply masked updates on a reconstruction network $g(\cdot;\phi)$ and a predictor model $f(\cdot;\theta)$ during training. During inference, the prediction will be generated directly by the updated $f(\cdot;\theta)$ with no additional cost. To further generalize across models of varying complexities, we also propose to apply Spectral Norm Regularization [snr, sngan] to parameters of the first or last linear layers without altering optimizers like previous works [samformer].
@@ 2.2 Proposed $g(\cdot;\phi)$ for Reconstruction
As depicted in Figure 2, the proposed $g(\cdot;\phi)$ comprises four convolutional layers, a cross-layer concatenation for multi-resolution integration, and a lightweight feedforward network (FFN) to decode the reconstructed results. The convolutional layers primarily act as a parameterized smoothing mechanism, similar to techniques for seasonal-trend decomposition [dlinear, sparsetsf, cyclenet]. The FFN then mixes information from different positions in a time series to reconstruct each data point using features extracted by convolutional layers (further details in Appendix D). Without introducing extra noise or patch-wise/point-wise masks [patchtst], we directly learn an identity mapping.
In essence, $g(\cdot;\phi)$ is designed to traverse diverse alternatives with varying levels of fidelity, which differs in purpose from the usual reconstruction task [ssl_con_gen, ssldenoisefinance, timesurl]. Due to this difference, the architecture of $g(\cdot;\phi)$ might benefit from novel designs; however, this is not examined in this study. The proposed $g(\cdot;\phi)$ is merely a simple prototype to validate our ideas and is not claimed superior to other unexplored options for this novel task.
@@ 3.3 Self-Correction with Adaptive Mask (Scam)
Further evidence is provided by analyzing the loss landscape. A sharpness metric for optimization, proposed by [samformer], measures the generalization capability of a model. Specifically, the sharpness metric is defined as $\lambda_{max}=\|\nabla^{2}_{\theta}\mathcal{L}\|_{2}^{2}$, the largest eigenvalue of the Hessian matrix. A higher $\lambda_{max}$ indicates a sharper loss landscape, which correlates with a worse generalization or more severe overfitting. By computing this metric on the converged parameters, we observe that $\ell_{target}\odot M$ does exhibit a sharper landscape compared to $\ell_{target}\odot\overline{M}$. To address this, we keep the less sharp part of the $\mathcal{L}_{sup}$ term, i.e., $\mathcal{L}_{sup}\odot\overline{M}$, resulting in the revised loss:
@@ 3.4 Spectral Norm Regularization (snr)
Note that we have omitted the gradient constraints in Eq. 3. This term is, in fact, positively correlated with $\nabla_{\theta}f(x;\theta)$ (discussed in Appendix E). This further supports our previous analysis, as we have only tested MLP models, which converge easily with lower $\nabla_{\theta}f$. However, when the predictor is replaced by a Transformer-based model, the dominant source of overfitting shifts from $\ell_{target}$ to $\ell_{pred}$. This phenomenon is not unique to our self-supervised learning paradigm. For example, [samformer] proposed sharpness-aware optimization to address overfitting in supervised settings. While gradient penalties are theoretically effective, they may not be practical for complex models like Transformers, as computing second-order derivatives can significantly hinder optimization.
@@
A more direct approach is to regularize parameters using the sharpness metric, known as Spectral Norm Regularization [sngan, snr]:
@@
where $\|\cdot\|_{2}$ is the spectral norm (the largest eigenvalue of parameter matrix) and $\gamma$ is a learnable scale factor. When applying to self-attention (SA) in Transformer-based architecture, snr significantly undermines the expressiveness of attention score matrices (entropy collapse [collapse]). Hence, samformer conclude that snr is inapplicable to SA parameters. However, we observe that linear layers — typically the embedding layer before SA and the projection layer after SA — also contribute to the overall sharpness of the loss landscape. Consequently, we propose applying snr selectively to the pre- and post-SA linear layers. In this way, snr can work with MLP-based models as they are also composed of multiple linear layers. Further empirical studies on snr are discussed in Section E.3.
@@ 4 Experiment and Analysis
Our main evaluation involves seven datasets: Electricity, Weather, Traffic, and four ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), all of which are well-established TSF benchmarks and publicly available [autoformer]. We also test the proposals using four PeMS datasets of a larger scale, as reported in Appendix 7. The predictor (i.e., the backbone model integrated with Scam) covers representative TSF models, including both MLP-based and Transformer-based architectures. Mlp [rlinear] is a vanilla 2-layer baseline equipped with RevIN [revin] while cycleNet [cyclenet] is a SOTA MLP-based model explicitly capturing cyclic trend in time series. PatchTST [patchtst] and iTransformer [itrans] are Transformer-based models, representing channel-independent and channel-dependent methods, respectively. Following previous settings [informer, autoformer, timesnet] for direct and fair comparison, we set prediction length $H\in\{96,192,336,720\}$ and look-back length to 96 for all datasets. We provide dataset descriptions, implementation details, and reproduction instructions in Appendix G.
@@ 4.1 Main Experiment (Q1)
Regarding Q1, our method demonstrates general effectiveness across various backbones and datasets. A well-known discrepancy between MLP-based and Transformer-based models is their dataset preferences: Transformer-based methods excel on large, regular datasets, while MLP-based methods perform better on noisy datasets [tfb, benchmark_tkde]. Scam helps bridge this gap by enabling Transformer-based models to perform competitively on traditionally challenging datasets and enhancing the robustness of MLP-based models.
@@ 4.2 Ablation Study (Q2)
snr, a practical alternative to the gradient penalty in Eq. 8, consistently enhances performance across backbones. However, iTrans, being more prone to overfitting on small datasets (discussed in [samformer]), benefits less from snr compared to cycleNet, likely due to insufficient data for generalization in higher-complexity architectures. With Scam, both models exhibit significant improvements, leveraging the expanded effective training data.
@@ 4.3 Scam: A Multiple Instance Learning View (Q3)
Multiple Instance Learning (MIL [MIL, milsurvey]) is a classical weakly-supervised binary classification problem where typically a bag of instances is labeled positive if at least one instance is positive. [millet, timemil] extend MIL to Time Series Classification by treating an input window as a bag of instances (time points), enabling a model to predict based on instance-level classifications that are more interpretable.
@@ 5 Related Work
Meta-Learning for Time Series. Meta-Learning, by definition, seeks to perform in a learn-to-learn paradigm. Generally speaking, meta-learning for TSF includes optimization-based and metric-based methods. Optimization-based methods target optimal initial parameters [maml], often involving a two-loop optimization process [metats, deeptime]. Metric-based methods [adarnn, deeptime] learn a metric function that provides more expressive measurements of distances between data samples, commonly comparing training and validation samples.
Our method Scam aligns with the broader scope of meta-learning. Specifically, the grid search (Section 3.1, Appendix B) follows the two-loop structure similar to optimization-based methods. However, it diverges by focusing on dataset space rather than parameter space. Additionally, the final mask form of Scam in essence provides a more accurate metric tailored for a supervised-learning setting. Previous works [deeptime, adarnn] learn metrics on the sample level (a window of time series), whereas ours focuses on the instance level (individual data points of time series).
Self-supervised Learning for Time Series. Self-supervised learning trains models without relying on manually labeled data by using auxiliary tasks like generation, contrast, and reconstruction to learn expressive representation or create pseudo labels. In the realm of time series, this approach is discussed more in Time Series Classification (TSC) [sslclassification1, sslclassification2, ssl_con_gen]. Recent works [millet, timemil] present a novel perspective that instances in time series/segments can have multiple labels. They propose corresponding weakly-supervised-learning methods that significantly improve both performance and interpretability.
@@ Appendix A Limitations
Our work constrains the discussion for TSF models in MLP- and Transformer-based ones, excluding other potentially feasible solutions [moderntcn, timemachine]. We omit these newly explored architectures since they have not outperformed the baseline methods we have adopted under the same budget. Also, we have not verified our method on LLM-based models (or models of an LLM-level scale) [moirai, chronos] due to the excessive costs required.
@@ Appendix C Further Discussion on Gradient Constraint
For readers who are familiar with Reinforcement Learning, this derivation resembles the transfer from a constrained optimization to a penalized one (e.g., from TRPO [trpo] to PPO [ppo]). In brief, while the penalized form is theoretically equivalent to the constrained form, it is challenging to choose a fixed $\beta$ that works universally across all datasets or even within a single dataset (because intrinsic characteristics can vary over time). Thus, a more general form of constraint is required to better serve the penalty, similar to the concept of gradient clipping in PPO.
@@
However, in practice, snr have limitations when applied to the parameter matrix in self-attention mechanisms. This phenomenon is termed entropy catastrophe as discussed by [samformer]. In this paper, by analyzing the sharpness of different components in the predictor model, we propose to use pre-snr and post-snr combined, which specifically normalizes the first and last linear layer in the TSF models (see Section 3.4).
@@ D.1 Conv-concat Layer
Effective Receptive Field. As what was proposed by [moderntcn], the Effective Receptive Field (ERF) is a reasonable consideration for designing convolution-based architectures. To evaluate the ERF of our conv-concat layer, we input an impulse function and visualize the resulting ERF, as shown in Figure 10. The visualization demonstrates that, without requiring an extra-large convolution kernel, our proposed method achieves a near-global ERF. This is made possible by combining the outputs from different layers, each capturing distinct frequency patterns.
@@ D.2 Point-wise FFN
  • Patch-dependent Design: PatchTST [patchtst] adopts a patch-dependent linear projector (shown in Figure 11(a)) that first flattens all features within patches and utilizes an extra-large weight matrix of shape $Ld\times d$, where $L$ is the sequence length and $d$ is the dimension of the latent embedding.
  • Patch-independent Design: [patchindependent] proposes that the necessity for patch-dependent designs depends on the specific task. For instance, tasks like forecasting may require patch-dependent projectors, while tasks like contrastive learning might favor a patch-independent design (shown in Figure 11(b)).
@@ E.2 Demystifying the Self-supervision in Scam
Channel-wise Distribution Alignment. Distribution shift, a core challenge in long-term Time Series Forecasting (TSF), is decisive for a TSF model to generalize on future data after training. The most widely adopted approach to address this issue is Reversible Instance Normalization (RevIn) [revin, nonstationary], which aligns the distributions of historical and future data.
@@
For models designed with channel-independence (CI), such as PatchTST [patchtst] and cycleNet [cyclenet], aligning distributions across channels is of critical importance, especially when trained on datasets with a large number of variates. The channel-wise alignment introduced by our self-supervised reconstruction task provides an effective solution to distribution shifts, thereby enhancing the performance of CI predictor models.
@@ E.3 In-depth Examination on snr
We further verify the effectiveness of snr by analyzing the loss landscape. Transformer-based models, as discussed, are more prone to overfitting, particularly on noisy datasets. This weakens our adaptive mask’s ability to discard overfitted components, as $\ell_{pred}$ — the predictor’s training loss — also contributes to overfitting. Gradient penalty offers a solution to this issue. For TSF models specifically, where both inputs and outputs are time series with abrupt distribution shifts, parameter robustness is critical for mitigating overfitting. Previous works [snr, sngan] have shown that regulating parameters via their spectral norm enhances stability against input perturbations, which is particularly beneficial in TSF.